# Shanghai 200433, China

### v1

@@ 3 Free Field Computation
Note also that in Equation (10), the coefficient of the logarithmic term is $-\frac{1}{2}$, which is different from $-1$ in this case. This is because that result is for compact scalar, whereas here we consider a non-compact scalar.
We can further analyze the $q$-dependence of $S(q)$. The expression (48) shows that $S(q)$ remains independent of the conserved charge $q$ up to the constant term. This reproduces the equipartition property. To explicitly show the $q$-dependence, we expand the expression up to $\mathcal{O}(V^{-2})$. This term depends on $q$ through a $q^{2}$ dependence, which is reasonable given that the complex scalar field is symmetric under charge conjugation. This also term sets a constraint on the validity of the large $V$ approximation, which gives $q\ll\log\left(\frac{R}{\delta}\right)$ in 2D scalar case.
In summary, for a conformal free complex scalar field in 2D, the expansion of symmetry-resolved entanglement entropy $S(q)$ in terms of $V$ consists of the conventional entanglement entropy $S$, plus the subleading term $-\log V$, followed by further subleading terms. The equipartition property holds up to the $\mathcal{O}(V)$ term, and breaks down at the next higher-order term. In Section 5, we will show that this is a consequence of the expansion in $V$, and that the dependence on $q$ can only appear in terms beyond the constant order.
@@
$\mathcal{Z}_{n}(q)=\int_{0}^{2\pi}\frac{\mathrm{d}\mu}{2\pi}\mathrm{e}^{V(g_{n}(\mu)-\mathrm{i}\mu\sigma)}.$ (53)
@@ 4 Holographic Computation
In this section, we calculate symmetry-resolved entanglement entropy in higher-dimensional holographic theories. We still employ the conformal mapping method introduced in Section 2 to transform the density matrix $\rho_{\mathcal{D}}$ of a spherical region into a thermal density matrix in hyperbolic space, as shown in Equation (12). Through the AdS/CFT correspondence Maldacena:AdSCFT ; Witten:AdS_and_holography ; GKP , the thermal state in hyperbolic space is further dual to a black hole in the bulk theory CHM:towards . Following the interpretation of $\mathrm{e}^{\mathrm{i}\mu Q}$ discussed in Section 2, we need to consider charged black holes in the bulk gravity theory, and $Z(n\beta_{0},\mathrm{i}\mu/(n\beta_{0}))$ in (16) is the grand-canonical partition function with imaginary chemical potential of the dual charged black hole. Using the classical gravity saddle-point approximation, we approximate the grand-canonical partition function with the on-shell action satisfying the required conditions. After computing the desired charged moments and performing the Fourier transforms, we obtain the symmetry-resolved Rényi entropy and symmetry-resolved entanglement entropy.
@@
The black hole mass [^4] The relationship between parameter $m$ and black hole mass $M$ is $M=\frac{R}{L}(d-1)V_{\Sigma}/({2l_{\mathrm{P}}^{d-1}})$. is related to the horizon radius $r_{\mathrm{h}}$ through
@@
The next-order term is the logarithmic contribution $-\frac{1}{2}\log V$, which introduces new universal terms beyond those present in the original entanglement entropy. For odd $d$, the universal term in $S(q)$ takes the form
@@ 5.2 Expansion structure via asymptotic analysis
$\mathcal{Z}_{n}(q)=\int_{0}^{2\pi}\frac{\mathrm{d}\mu}{2\pi}\mathrm{e}^{-\mathrm{i}\mu q}\mathcal{Z}_{n}(\mu)=\int_{0}^{2\pi}\frac{\mathrm{d}\mu}{2\pi}\mathrm{e}^{-\mathrm{i}\mu q}\mathrm{e}^{Vg_{n}(\mu)}.$ (85)
In the asymptotic analysis for this integral transform, both functions $\mathrm{e}^{-\mathrm{i}\mu q}$ and $\mathrm{e}^{Vg_{n}(\mu)}$ need to be expanded around the maximum point of $g_{n}(\mu)$ at $\mu=c$ (specifically, $c=0$ in the calculations of the previous chapter) Bender:asymp . The leading behavior of function $g_{n}(\mu)$ near its maximum is controlled by its first few derivatives. If $g^{\prime}_{n}(\mu)=0$ at the maximum point $\mu=0$ (for 4D scalar in Section 3.1 and holographic theories in Section 4), then beyond the function value $g_{n}(0)$ itself, its leading behavior is controlled by the second derivative; if $g^{\prime}_{n}(\mu)\neq 0$ (for 2D scalar in Section 3.2), then after the function value, the next most important term is the first derivative. Below we examine both cases.
@@
$\mathcal{Z}_{n}(q)=\frac{\mathrm{e}^{Vg}}{2\pi}\frac{1}{\sqrt{V(-g^{\prime\prime})}}\left(h+\frac{-\frac{h^{\prime\prime}}{2g^{\prime\prime}}+\frac{g^{(3)}h^{\prime}}{2(g^{\prime\prime})^{2}}-\frac{5h\left(g^{(3)}\right)^{2}}{24(g^{\prime\prime})^{3}}+\frac{hg^{(4)}}{8(g^{\prime\prime})^{2}}}{V}\right).$ (86)
Here, $h(\mu)=\mathrm{e}^{-\mathrm{i}\mu q}$, $g^{(k)}$ denotes the $k$-th derivative with respect to $\mu$, and both the function $h$ and the derivatives of $g$ are evaluated at $\mu=c$ ($c=0$ for the cases we consider in this paper). The reason that it’s $\sqrt{V}$ in the denominator of $\mathcal{Z}_{n}(q)$ is because the second-derivative term is the primary term after the function value. The derivation and explanation of this result are discussed in Appendix B.
@@
$\mathcal{Z}_{n}(q)=\frac{\mathrm{e}^{Vg_{n}(c)}}{2\pi\sqrt{Vf_{n}}}\left(1+\frac{a_{n}(q)}{V}\right),$ (87)
@@
$S_{n}(q)=V\frac{(ng_{1}(c)-g_{n}(c))}{n-1}-\frac{1}{2}\log V\\ +\frac{-n\log f_{1}+\log f_{n}}{2(n-1)}-\log(2\pi)+\frac{1}{V}\frac{na_{1}(q)-a_{n}(q)}{n-1}.$ (88)
@@
$S(q)=V\left(g_{1}(c)-\left.\frac{\partial g_{n}(\mu)}{\partial n}\right|_{n=1,\mu=c}\right)-\frac{1}{2}\log V+\frac{f^{\prime}(1)}{2f(1)}-\frac{1}{2}\log(f(1))-\log(2\pi)\\ +\frac{1}{V}\left(a_{1}(q)-\left.\frac{\partial a_{n}(q)}{\partial n}\right|_{n=1}\right).$ (89)
@@
In this case, the first-derivative term is the most important term after the function value at $\mu=0$. The general expression for the Fourier integral result is approximately given by
@@
This conforms to the expansion structure described in Section 5.1—the first term is the unresolved entanglement entropy; the second term is $-\log V$; the next term is constant, and the SREE remains independent of $q$ up to this order, exhibiting the equipartition property; the breaking of equipartition appears at order $V^{-1}$. Note that $a_{n}=2n$ in 2D scalar in Section 3.2, thus the coefficient of order $(V^{-1})$, i.e. $a_{1}-\left.\partial_{n}a_{n}\right|_{n=1}$, vanishes in that case.
@@ A.1 Entropy and the coarse-graining property
Figure 3: Illustration for microstates (big black dots) and sub-microstates (small dots in small circle). The “microstate” $i$ appears with probability $p_{i}$ and has its own entropy $S_{i}$. The actual total entropy comes from uncertainty of probability distribution $\{p_{i}\}$ and uncertainty within state $i$.
@@
Given Shannon’s definition of entropy, this formula can be proved. However, it’s worth noting that Shannon, taking the opposite route, took (95) as one of the defining properties to derive his expression of entropy Shannon . Note that the argument of total entropy $S$ now is more detailed probability distribution, which we don’t explicitly write down. Also note that, at a deeper level, there might be even more uncertainty in $S_{i}$, and $S_{i}$ can be decomposed in the same way as $S$ in (95). In this sense, $S_{i}$ and $S$ are in fact of the same nature. This will still be true in quantum physics, where both $S_{i}$ and $S$ are entanglement entropy as discussed below in the next subsection, while $-\sum_{i}p_{i}\log p_{i}$ is still classical Shannon entropy.
@@ A.2 In quantum theory: symmetry-resolved entanglement
Consider a bipartite system with global U(1) symmetry with conserved charge $Q$, which is a sum of charge of the two subsystems, $Q=Q_{A}+Q_{B}$. When the total system density matrix $\rho$ commutes with the conserved charge $Q$, it can be proved that, by partial trace, the subsystem density matrices still commute with the corresponding charge operator, i.e., $[\rho_{A},Q_{A}]=[\rho_{B},Q_{B}]=0$. Therefore, the subsystem density matrix $\rho_{A}$ (let’s focus on subsystem $A$ from now on) will be block-diagonal in the eigenbases of the conserved charge $Q_{A}$:
@@
$S(\rho_{A})=-\sum_{i}p(q)\log p(q)+\sum_{q}p(q)S(\rho_{A}(q)).$ (102)
Appendix B Asymptotic Analysis of Fourier Integral

### v2

@@ 3 Free Field Computation
Note also that in Equation (10), the coefficient of the logarithmic term is $\frac{1}{2}$, which is different from 1-1 in this case. This is because that result is for compact scalar [^5] Due to its twisted boundary condition, a compact scalar with identification $\phi\sim\phi+2\pi$ has a different heat kernel than that of a non-compact real scalar. This lies beyond the scope of the present work, in which we discuss only non-compact scalars., whereas here we consider a non-compact scalar. In DiGiulio:BCFT_to_SRE , the authors also analyze a non-compact (real) scalar in 2D, focusing on the translation symmetry $\phi\rightarrow\phi+a$; by contrast, we focus on the U(1) rotational symmetry of the complex scalar in this work.
We can further analyze the $q$-dependence of $S(q)$. The expression (48) shows that $S(q)$ remains independent of the conserved charge $q$ up to the constant term. This reproduces the equipartition property. To explicitly show the $q$-dependence, we expand the expression up to $\mathcal{O}(V^{-2})$. This term depends on $q$ through a $q^{2}$ dependence, which is reasonable given that the complex scalar field is symmetric under charge conjugation. This term also sets a constraint on the validity of the large $V$ approximation, which gives $q\ll\log\left(\frac{R}{\delta}\right)$ in 2D scalar case.
In summary, for a conformal free complex scalar field in 2D, the expansion of symmetry-resolved entanglement entropy $S(q)$ in terms of $V$ consists of the conventional entanglement entropy $S$, plus the subleading term $\log V$, followed by further subleading terms. The equipartition property holds up to the $\mathcal{O}(V)$ term, and breaks down at the next higher-order term. In Section 5, we will show that this is a consequence of the expansion in $V$, and that the dependence on $q$ can only appear in terms beyond the constant order.
@@
$\mathcal{Z}_{n}(q)=\int_{-\pi}^{\pi}\frac{\mathrm{d}\mu}{2\pi}\mathrm{e}^{V(g_{n}(\mu)-\mathrm{i}\mu\sigma)}.$ (53)
@@ 4 Holographic Computation
In this section, we calculate symmetry-resolved entanglement entropy in higher-dimensional holographic theories. We still employ the conformal mapping method introduced in Section 2 to transform the density matrix $\rho_{\mathcal{D}}$ of a spherical region into a thermal density matrix in hyperbolic space, as shown in Equation (12). Through the AdS/CFT correspondence Maldacena:AdSCFT ; Witten:AdS_and_holography ; GKP , the thermal state in hyperbolic space is further dual to a black hole in the bulk theory CHM:towards . Following the interpretation of $\mathrm{e}^{\mathrm{i}\mu Q}$ discussed in Section 2, we need to consider charged black holes in the bulk gravity theory, and $Z(n\beta_{0},\mathrm{i}\mu/(n\beta_{0}))$ in Equation (16) is the grand-canonical partition function with imaginary chemical potential of the dual charged black hole. Using the classical gravity saddle-point approximation, we approximate the grand-canonical partition function with the on-shell action satisfying the required conditions. After computing the desired charged moments and performing the Fourier transforms, we obtain the symmetry-resolved Rényi entropy and symmetry-resolved entanglement entropy [^6] We note that charged moments in holographic settings have been investigated using similar techniques in Baiguera:shape_deform_charged_Renyi . They employed the conformal transformation and a dual black hole solution. However, there is a key difference: they examine a quantity called charged Rényi entropy proposed in Belin:holo_charged_renyi , which is a function of $\mu$, whereas we consider symmetry-resolved entropy, which is a function of charge $q$ and quantifies entanglement within charge sectors..
@@
The black hole mass [^7] The relationship between parameter $m$ and black hole mass $M$ is $M=\frac{R}{L}m(d-1)V_{\Sigma}/({2l_{\mathrm{P}}^{d-1}})$. is related to the horizon radius $r_{\mathrm{h}}$ through
@@
The next-order term is the logarithmic contribution $\frac{1}{2}\log V$, which introduces new universal terms beyond those present in the original entanglement entropy. For odd $d$, the universal term in $S(q)$ takes the form
@@ 5.2 Expansion structure via asymptotic analysis
$\mathcal{Z}_{n}(q)=\int_{-\pi}^{\pi}\frac{\mathrm{d}\mu}{2\pi}\mathrm{e}^{-\mathrm{i}\mu q}\mathcal{Z}_{n}(\mu)=\int_{-\pi}^{\pi}\frac{\mathrm{d}\mu}{2\pi}\mathrm{e}^{-\mathrm{i}\mu q}\mathrm{e}^{Vg_{n}(\mu)}.$ (85)
In the asymptotic analysis for this integral transform, both functions $\mathrm{e}^{-\mathrm{i}\mu q}$ and $\mathrm{e}^{Vg_{n}(\mu)}$ need to be expanded around the maximum point of $g_{n}(\mu)$ at $\mu=c$ (specifically, $c=0$ in the calculations of the previous section) Bender:asymp . The leading behavior of the function $g_{n}(\mu)$ near its maximum is controlled by its first few derivatives. If $g^{\prime}_{n}(\mu)=0$ at the maximum point $\mu=0$ (for 4D scalar in Section 3.1 and holographic theories in Section 4), then beyond the function value $g_{n}(0)$ itself, its leading behavior is controlled by the second derivative; if $g^{\prime}_{n}(\mu)\neq 0$ (as for 2D scalar in Section 3.2, where $g^{\prime}_{n}(\mu)$ is in fact discontinuous at $\mu=0$), then after the function value, the next most important term is the first derivative. Below we examine both cases.
@@
$\mathcal{Z}_{n}(q)=\frac{\mathrm{e}^{Vg}}{\sqrt{2\pi}}\frac{1}{\sqrt{V(-g^{\prime\prime})}}\left(h+\frac{-\frac{h^{\prime\prime}}{2g^{\prime\prime}}+\frac{g^{(3)}h^{\prime}}{2(g^{\prime\prime})^{2}}-\frac{5h\left(g^{(3)}\right)^{2}}{24(g^{\prime\prime})^{3}}+\frac{hg^{(4)}}{8(g^{\prime\prime})^{2}}}{V}\right).$ (86)
Here, $h(\mu)=\mathrm{e}^{-\mathrm{i}\mu q}$, $g^{(k)}$ denotes the $k$-th derivative with respect to $\mu$, and both the function $h$ and the derivatives of $g$ are evaluated at $\mu=c$ ($c=0$ for the cases we consider in this paper). The reason that it’s $\sqrt{V}$ in the denominator of $\mathcal{Z}_{n}(q)$ is because the second-derivative term is the primary term after the function value. The derivation and explanation of this result are discussed in Appendix C.
@@
$\mathcal{Z}_{n}(q)=\frac{\mathrm{e}^{Vg_{n}(c)}}{\sqrt{2\pi}\sqrt{Vf_{n}}}\left(1+\frac{a_{n}(q)}{V}\right),$ (87)
@@
$S_{n}(q)=V\frac{(ng_{1}(c)-g_{n}(c))}{n-1}-\frac{1}{2}\log V\\ +\frac{-n\log f_{1}+\log f_{n}}{2(n-1)}-\frac{1}{2}\log(2\pi)+\frac{1}{V}\frac{na_{1}(q)-a_{n}(q)}{n-1}.$ (88)
@@
$S(q)=V\left(g_{1}(c)-\left.\frac{\partial g_{n}(\mu)}{\partial n}\right|_{n=1,\mu=c}\right)-\frac{1}{2}\log V+\frac{f^{\prime}(1)}{2f(1)}-\frac{1}{2}\log(f(1))-\frac{1}{2}\log(2\pi)\\ +\frac{1}{V}\left(a_{1}(q)-\left.\frac{\partial a_{n}(q)}{\partial n}\right|_{n=1}\right).$ (89)
@@
In this case, the first-derivative term is the most important term after the function value at $\mu=0$. The general expression for the Fourier integral result is approximately given by [^11] The derivative is actually discontinuous at $\mu=0$. However, as explained in the previous footnote, to perform this asymptotic expansion we treat the cases $\mu=0^{+}$ and $\mu=0^{-}$ separately and then sum the results. the discontinuity presents no new conceptual difficulties.
@@
This conforms to the expansion structure described in Section 5.1—the first term is the unresolved entanglement entropy; the second term is $\log V$; the next term is constant, and the SREE remains independent of $q$ up to this order, exhibiting the equipartition property; the breaking of equipartition appears at order $V^{-1}$. Note that $a_{n}=2n$ in 2D scalar in Section 3.2, thus the coefficient of order $(V^{-1})$, i.e. $a_{1}-\left.\partial_{n}a_{n}\right|_{n=1}$, vanishes in that case.
@@ A.1 Entropy and the coarse-graining property
Figure 3: Illustration for microstates (big black dots) and sub-microstates (small dots in small circle). The “microstate” $i$ appears with probability $p_{i}$ and has its own entropy $S_{i}$. The actual total entropy comes from uncertainty of probability distribution $\{p_{i}\}$ and uncertainty within state ii.
@@
Given Shannon’s definition of entropy, this formula can be proved. However, it’s worth noting that Shannon, taking the opposite route, took (95) as one of the defining properties to derive his expression of entropy Shannon . Note that the argument of total entropy $S$ now is more detailed probability distribution, which we don’t explicitly write down. Also note that, at a deeper level, there might be even more uncertainty in $S_{i}$, and $S_{i}$ can be decomposed in the same way as $S$ in (95). In this sense, $S_{i}$ and $S$ are in fact of the same nature. This will still be true in quantum physics, where both $S_{i}$ and $S$ are entanglement entropy as discussed below in the next subsection, while $\sum_{i}p_{i}\log p_{i}$ is still classical Shannon entropy.
@@ A.2 In quantum theory: symmetry-resolved entanglement
Consider a bipartite system with global U(1) symmetry with conserved charge $Q$, which is a sum of charge of the two subsystems, $Q=Q_{A}+Q_{B}$. When the total system density matrix $\rho$ commutes with the conserved charge $Q$, it can be proved that, by partial trace, the subsystem density matrices still commute with the corresponding charge operator, i.e., $\rho_{A},Q_{A}]=[\rho_{B},Q_{B}]=0$. Therefore, the subsystem density matrix $\rho_{A}$ (let’s focus on subsystem $A$ from now on) will be block-diagonal in the eigenbases of the conserved charge $Q_{A}$:
@@
$S(\rho_{A})=-\sum_{q}p_{A}(q)\log p_{A}(q)+\sum_{q}p_{A}(q)S(\rho_{A}(q)).$ (102)
Appendix B Heat Kernel Essentials
In this section, we present a concise, pedagogical introduction to the essentials of heat kernel in QFT needed for the main text, and derive key expressions, such as the heat kernel with twisted boundary condition on $S^{1}$ (Equation (28)).
B.1 Basic definition of heat kernel
The heat kernel $K(x,y,t)$ is originally defined as the fundamental solution to the heat equation:
$\left(\partial_{t}+D_{x}\right)K(x,y,t)=0,$ (103)
with the initial condition as a point source
$K(x,y,0)=\delta(x-y),$ (104)
where $D_{x}$ is a differential operator (often the Laplacian or Laplace-type operator) acting on the $x$ variable.
In QFT, the heat kernel is a powerful tool for evaluating functional determinants and partition functions. As an example of heat kernel in QFT, we consider the (Euclidean) path integral representation of the generating functional for real scalar field $\phi$:
$Z=\int\mathcal{D}\phi\,\exp(-S[\phi])$ (105)
where the action $S$ only contains the quadratic part (free field):
$S=\frac{1}{2}\int\phi D\phi\sqrt{g}\mathrm{d}^{d}x$ (106)
Here, $D$ is the differential operator
$D=D_{0}:=-\nabla^{\mu}\nabla_{\mu}.$ (107)
The heat kernel can be expressed as:
$K_{D}(x,y,t):=\langle x|\mathrm{e}^{-tD}|y\rangle.$ (108)
One can easily verify that $K(x,y,t)$ satisfies the heat equation with the initial condition being a $\delta$ function source:
$\begin{cases}(\partial_{t}+D_{x})K_{D}(x,y,t)=0,\\ K_{D}(x,y,0)=\delta(x,y).\end{cases}$ (109)
The heat kernel solution to this equation depends on the differential operator $D$ and the boundary condition.
For the complex scalar we consider in Section 3, the path integral can be evaluated by a Gaussian integral, yielding
$Z\sim\frac{1}{\det D}.$ (110)
The free energy $F=-\log Z$ can be computed by Vassilevich:heatkernel_manual
$F=-\log Z=\log\det D=-\int_{0}^{\infty}\frac{\mathrm{d}t}{t}\,\operatorname{tr}\left[\mathrm{e}^{-tD}\right].$ (111)
Using expression (108), we can finally write $F$ in terms of the heat kernel:
$F=-\int_{0}^{\infty}\frac{\mathrm{d}t}{t}\int\mathrm{d}^{n}x\sqrt{g}K_{D}(x,x,t).$ (112)
Therefore, the computation of the generating functional (and charged moments needed in this paper) for the free field can be completed using heat kernel techniques. The specific form of the heat kernel depends on the differential operator $D$ and any imposed boundary condition (in this paper, twisted boundary conditions on the Euclidean time circle).
We also note the connection to the real (non-compact) scalar. A free complex scalar field can be viewed as two independent real scalars, so its action decomposes into the sum of two identical real-scalar actions. Evaluating again the Gaussian integral yields partition function $Z_{\text{real}}\sim\frac{1}{\sqrt{\det D}}$. Hence, for $\mu=0$, the free energy of complex scalar is exactly twice that of that of real scalar: $F_{\text{complex}}=2F_{\text{real}}$.
The above derivation of the heat kernel is rather schematic; for a more rigorous mathematical treatment and its applications in QFT, see, for example, Vassilevich:heatkernel_manual .
B.2 Heat Kernel on $S^{1}_{\beta}$ with twisted boundary condition
Here we derive the heat kernel of the Laplace operator on $S^{1}$ in Equation (28) with twisted boundary condition (25).
Starting from the heat kernel as a “transition amplitude” (108), by inserting an orthonormal basis in it, we obtain the heat kernel as a “mode sum”:
$K(x,y,t)=\sum_{n}\Braket{x|\phi_{n}}\Braket{\phi_{n}|\mathrm{e}^{-tD}|y}=\sum_{n}\phi_{n}(x)\phi_{n}^{*}(y)\mathrm{e}^{-t\lambda_{n}},$ (113)
where $\phi_{n}$ are eigenfunctions of $D$ with eigenvalues $\lambda_{n}$.
Consider a circle $S^{1}_{\beta}$ of circumference $\beta$, and a field $\phi(x)$ satisfying the twisted boundary condition:
$\phi(x+\beta)=\mathrm{e}^{\mathrm{i}\mu}\phi(x).$ (114)
The corresponding eigenfunctions and eigenvalues are
$\phi_{n}(x)=\frac{1}{\sqrt{\beta}}\mathrm{e}^{\mathrm{i}\frac{2\pi n+\mu}{\beta}x},\,\lambda_{n}=\left(\frac{2\pi n+\mu}{\beta}\right)^{2},\,n\in\mathbb{Z}.$ (115)
Substituting this into the mode sum and using the Poisson summation formula to convert it to an “image sum”, we obtain
$K_{S^{1}_{\beta}}(x_{1},x_{2},t)=\frac{1}{\sqrt{4\pi t}}\sum_{k\in\mathbb{Z}}\exp\left(-\frac{(x_{1}-x_{2}-\beta k)^{2}}{4t}+\mathrm{i}k\mu\right).$ (116)
By setting $\beta=2\pi n$, this expression recovers Equation (28).
Appendix C Asymptotic Analysis of Fourier Integral